Multi-period Inventory Control Question

The average daily demand for sewing needles at RD Designs is 300 units with a standard  deviation of 50 units. An order takes 2 days to arrive and the representative stops by every  6 days. The CEO of the RD Designs desires to maintain a 99% service level. Assuming the  representative just arrived, and there are 150 needles in stock, how many needles should  they order? 
Answer #1

To calculate the order quantity, we need to consider the lead time demand and safety stock to achieve the desired service level.

Given information: Average daily demand (D) = 300 units Standard deviation of daily demand (σ) = 50 units Lead time (LT) = 2 days Visit frequency (VF) = 6 days Service level = 99% (which corresponds to a z-value of 2.33 for a normal distribution)

Step 1: Calculate the Lead Time Demand (LTD) LTD = D * LT LTD = 300 units/day * 2 days = 600 units

Step 2: Calculate the Safety Stock (SS) using the z-value SS = z * σ * sqrt(VF) SS = 2.33 * 50 units * sqrt(6 days) ≈ 2.33 * 50 units * 2.449 = 286.285 ≈ 287 units

Step 3: Calculate the Reorder Point (ROP) ROP = LTD + SS ROP = 600 units + 287 units = 887 units

Step 4: Calculate the Order Quantity (Q) Q = ROP - Current Stock Q = 887 units - 150 units = 737 units

The company should order 737 needles to maintain a 99% service level with a representative just having arrived and 150 needles in stock.


answered by: mervetokaz